Sir, i am a student of class 9 and i have a doubt.


The question is.. From the top of a tower of 200m height, a stone is projected vertically upwards with a velocity of 49 m/s.


Calculate the time to reach the ground.


I Did approach the answer through a website .. but i dint understand why the initial velocity (u) was with a negative sign.


Sir.. can u show me the method and the explanations of my doubts,.. Please

Grade:12

1 Answers

Kurma Anudeep
11 years ago

let s=200m

u=-49m/s

a=9.8m/s^2

v^2=u^2+2as

      =2401+19.6*200

       =2401+3920

v=79.5m/s

v=u+at

79.5=-49+9.8t

128.5=9.8t

t=128.5/9.8=13.11s

u is negative because the ball is thrown against the gravitational force

ALL THE BEST